# Advaith Sethuraman, 2505 Hayward St., Ann Arbor, MI 48109

### v1

@@ 4.2.3 Terrain Type Assessment:
We select the terrain type as a category for determining the test-train split, as the varying conditions of terrain can play a crucial role in the detection and segmentation of shipwrecks. This is most prominent in terrains that have a large amount of texture, making it difficult even for humans to discern the shipwreck from the terrain, leading to either false negatives or false positives. We hence aim to have the terrains split evenly based on their characteristics observed from the SSS imagery. In order to determine the most even split across the data based on terrain, we utilize clustering methods.
Specifically, the clustering is done on the N-dimensional latent space of a Variational Auto Encoder (VAE) pre-trained on a large color image (RGB) dataset. The input images are pre-processed by applying white balancing. This is done in order to prevent the clustering based on image intensity; otherwise, the intensity differences across the dataset are too prominent. We use $k$-means clustering, yielding a partitioning on an image-to-image basis. Each image is grouped based on the site it belongs to, which is then used to inform the test-train split. We assign sites a class of A for terrains with small-scale texture, a class of B for terrains with large-scale texture, a class of C for smooth terrain, and a class of D for patchy terrain texture.
@@ 4.3 Dataset Statistics
The dataset consists of $286$ images exported from the Discover software as high-resolution images. Note that despite the number of images being low, the average dimension of the images is $3,480\times 1,728$ pixels. The width of each exported image is fixed at $1,728$ pixels. The height varies across the $286$ images due to it being dependent on the leg length of each survey. We note that by applying the strided-cropping, as described in Section 3.5, we can obtain $4,672$ images of shape $728\times 1,$ 728 × 1 , 728 pixels. Table 1 provides further details of each site, including material, ship length and ship beam (width) obtained from Thunder Bay National Marine Sanctuary (Accessed online: 2024), as well as classification details according to our data categorization for terrain type, sonar image quality, and wreck condition.
@@ 5 Experiments & Results
Table 2: Aggregate baseline performance averaged across sites: metrics are computed per site and then the average is taken across all test sites for each metric. Metrics include mIoU, per-class IoU metrics, F1 Score, TPR, TNR, FPR, and FNR. $\uparrow$ indicates higher is better. $\downarrow$ indicates lower is better.
Baseline
mIoU $\uparrow$
[…]
Yang et. al
0.609
0.229 0.990 0.333 0.319 0.996 0.004 0.681
ViT-Adapter
0.637
0.283 0.991 0.395 0.488 0.995 0.005 0.512
DeepLabv3
0.678
0.363 0.992 0.473 0.485 0.996 0.004 0.515
HRNet
0.683
0.372 0.994 0.490 0.516 0.998 0.002 0.484
UNet
0.702
0.411 0.994 0.526 0.592 0.997 0.003 0.408
SOD 0.720 0.445 0.994 0.594 0.652 0.997 0.003 0.348
Table 3: Per-site mIoU compared across each baseline for each site. Higher is better ($\uparrow$).
Artificial Reef
Barge No. 1
Corsair
Corsican
Davidson
W.H. Gilbert
Haltiner Barge
L. Van Valkenburg
Monohansett
Monrovia
Shamrock
W.P. Thew
Viator
Yang et. al
0.496
0.735 0.616 0.527 0.492 0.635 0.515 0.668 0.499 0.730 0.548 0.566 0.778
ViT-Adapter
0.498
0.704 0.771 0.541 0.499 0.501 0.680 0.564 0.736 0.724 0.673 0.576 0.812
DeepLabv3
0.495
0.834 0.765 0.518 0.514 0.828 0.730 0.814 0.515 0.693 0.572 0.700 0.831
HRNet
0.503
0.851 0.646 0.586 0.533 0.819 0.724 0.813 0.602 0.775 0.498 0.710 0.820
UNet
0.499
0.867 0.781 0.555 0.536 0.862 0.621 0.813 0.518 0.779 0.712 0.765 0.821
SOD 0.507 0.887 0.791 0.532 0.518 0.874 0.721 0.850 0.732 0.779 0.497 0.783 0.886
@@ 5.1 Baseline Comparison
We perform benchmark evaluations on a series of deep learning-based segmentation methods to evaluate the ability of state-of-the-art deep learning methods to segment shipwrecks from SSS imagery. Benchmark methods are selected to reflect the variety of deep learning-based segmentation models used in the vision community: Yang et al. Yang et al. (2022), ViT Adapter Chen et al. (2023), DeepLabV3 Chen et al. (2017), HRNet Wang et al. (2020), UNet Ronneberger et al. (2015), and Salient Object Detection (SOD) InSPyReNet Kim et al. (2022). The model with the longest training time was SOD, which took 14 hours on a single NVIDIA A100 GPU with 80Gb of memory. All baselines were trained on $512\times 512$ images.
Figure 9: Shipwreck segmentation predictions from the baselines on Barge No. 1, Monrovia, and Haltiner
Each model is evaluated on an aggregate level with mean Intersection Over Union (mIoU), IoU for the “shipwreck” class ($\text{IoU}_{\text{ship}}$), IoU for the “other” class ($\text{IoU}_{\text{other}}$), F1 score, True Positive Rate (TPR), True Negative Rate (TNR), False Positive Rate (FPR), and False Negative Rate (FNR). Table 2 provides results across baselines, averaged across each site. Table 3 shows mIoU per site for each baseline. Segmentation predictions for all six baselines on three example sites are shown in Fig. 9. Across this comparison, the SOD model consistently outperformed or performed comparably against all other networks for the aggregate metrics. We performed all subsequent experiments with the SOD model.
@@ 6 Conclusion & Future Work
This work contributes AI4Shipwrecks, an open-source dataset for comparison of state-of-the-art deep neural networks for shipwreck segmentation from sonar imagery. While recent advances in deep learning have revolutionized the field of computer vision for terrestrial robotics, adoption of similar methods across marine applications is limited due to a lack of widely available data for development and direct comparison of results. We establish a benchmark for semantic segmentation of shipwrecks on the AI4Shipwrecks dataset, and we provide comparison of current state-of-the-art deep neural networks for segmentation. The dataset and code for evaluation will be open-source to enable future research in machine learning for ocean exploration.
A promising direction for future work includes leveraging synthetic data to augment real sonar datasets for learning-based detection and segmentation tasks Lee et al. (2018); Sethuraman et al. (2022). Additionally, work should also focus on advancing network architecture to enable DNNs to learn from limited training data. Of notable interest, recent work has demonstrated the potential for few shot learning for object detection from marine optical and sonar imagery Ochal et al (2020). Few shot and one shot learning aim to effectively learn to represent a class of objects after seeing a few or single instance of that class, which is ideal for datasets with relatively low abundance of samples per class.
@@
We would like to acknowledge the lives that were lost in shipwrecks throughout Thunder Bay National Marine Sanctuary. We would also like to thank the sanctuary scientists at Thunder Bay National Marine Sanctuary for supporting field experiments and labeling efforts.

### v2

@@ 4.2.3 Terrain Type Assessment:
We select the terrain type as a category for determining the train-test split, as the varying conditions of terrain can play a crucial role in the detection and segmentation of shipwrecks. This is most prominent in terrains that have a large amount of texture, making it difficult even for humans to discern the shipwreck from the terrain, leading to either false negatives or false positives. We hence aim to have the terrains split evenly based on their characteristics observed from the SSS imagery. In order to determine the most even split across the data based on terrain, we utilize clustering methods.
Specifically, the clustering is done on the N-dimensional latent space of a Variational Auto Encoder (VAE) pre-trained on a large color image (RGB) dataset. The input images are pre-processed by applying white balancing. This is done in order to prevent the clustering based on image intensity; otherwise, the intensity differences across the dataset are too prominent. We use $k$-means clustering, yielding a partitioning on an image-to-image basis. Each image is grouped based on the site it belongs to, which is then used to inform the train-test split. We assign sites a class of A for terrains with small-scale texture, a class of B for terrains with large-scale texture, a class of C for smooth terrain, and a class of D for patchy terrain texture.
@@ 4.3 Dataset Statistics
The dataset consists of $286$ images exported from the Discover software as high-resolution .PNG images. Out of these $286$ images, $161$ contain shipwrecks. There are $1.49$e77{7}7 shipwreck pixels and $4.39$e1111{11}11 background pixels.
Table 5: Aggregate baseline performance averaged across sites: metrics are computed per site and then the average is taken across all test sites for each metric. Metrics include IoUship, F1 Score, TPR, and TNR. TPR and TNR are calculated assuming the shipwreck class is positive. $\uparrow$ indicates higher is better.
{tabu}
c| c c c c Baseline IoUship $\uparrow$ F1 Score $\uparrow$ TPR $\uparrow$ TNR $\uparrow$
Yang et. al 0.212 0.310 0.296 0.995
ViT-Adapter 0.283 0.395 0.488 0.995
DeepLabv3 0.363 0.473 0.485 0.996
HRNet 0.372 0.490 0.516 0.998
UNet 0.411 0.526 0.592 0.997
SOD 0.445 0.594 0.652 0.997
Note that despite the number of images being low, the average dimension of the images is $3480\times 1728$ pixels. The width of each exported image is fixed at $1728$ pixels. The height varies across the $286$ images due to it being dependent on the leg length of each survey. We note that by applying the strided-cropping, as described in Section 3.6, we can obtain $2539$ images in the train set and $1722$ images in the test set of size $1728\times 1728$ pixels. Table 4 provides further details of each site, including material, ship length and ship beam (width) obtained from Thunder Bay National Marine Sanctuary (2024), as well as classification details according to our data categorization for terrain type, sonar image quality, and wreck condition.
Figure 9: Shipwreck segmentation predictions from the baselines on Barge No. 1, Monrovia, and Haltiner. Zoomed in for better detail.
@@ 5.1 Baseline Comparison
We perform benchmark evaluations on a series of deep learning-based segmentation methods to evaluate the ability of state-of-the-art deep learning methods to segment shipwrecks from SSS imagery. Benchmark methods are selected to reflect the variety of deep learning-based segmentation models used in the vision community: Yang et al. (Yang, Cheng, et al., 2022), ViT Adapter (Chen et al., 2023), DeepLabV3 (Chen et al., 2017), HRNet (Wang et al., 2020), UNet (Ronneberger et al., 2015), and Salient Object Detection (SOD) InSPyReNet (Kim et al., 2022). The model with the longest training time was SOD, which took 14 hours on a single NVIDIA A100 GPU with 80GB of memory. All baselines were trained on $512\times 512$ images, which are downsampled from the original square cropped size of $1728\times 1728$ described in Section 3.6. Code to replicate our cropping process and to set up a PyTorch dataset can be found in the repository linked on the project website.
We train each baseline according to the training parameters (learning rate, epoch, batch size) suggested by their original papers. This includes the choice of encoder backbones, for which we use the pre-trained backbone associated with the highest performance reported in the original papers. Although we use pre-trained backbones, we train each model from scratch on our dataset. We use the code released from the baseline papers, except for Yang et al., which we re-implemented ourselves following the architecture discussed in (Yang, Cheng, et al., 2022).
5.2 Evaluation Metrics
Each model is evaluated on an aggregate level with Intersection over Union (IoU) for the “shipwreck" class ($\text{IoU}_{\text{ship}}$), F1 score, True Positive Rate (TPR), and True Negative Rate (TNR). All metrics are calculated in pixel-space, and are therefore measures of how well network output corresponds to expert opinions in pixel-space. Table 4.3 provides results across baselines, averaged across each site. Table 4.2.1 shows $\text{IoU}_{\text{ship}}$ per site for each baseline. Segmentation predictions for all six baselines on three example sites are shown in Fig. 9. Across this comparison, the SOD model consistently outperformed or performed comparably against all other networks for the aggregate metrics. We performed the subsequent experiment with the SOD model.
246810120.250.250.250.250.30.30.30.30.350.350.350.350.40.40.40.40.450.450.450.45Number of Sites in Training SetAverage $\text{IoU}_{\text{ship}}$verage $\text{IoU}_{\text{ship}}$ vs. Number of Sites in Training Set
@@ 6 Conclusion & Future Work
This work contributes AI4Shipwrecks, an open-source dataset for comparison of state-of-the-art deep neural networks for shipwreck segmentation from SSS imagery. While recent advances in deep learning have revolutionized the field of computer vision for terrestrial robotics, adoption of similar methods across marine applications is limited due to a lack of widely available data for development and direct comparison of results. We establish a benchmark for semantic segmentation of shipwrecks on the AI4Shipwrecks dataset, and we provide comparison of current state-of-the-art deep neural networks for segmentation. The dataset and code for evaluation will be open-source to enable future research in machine learning for ocean exploration.
The AI4Shipwrecks dataset is a powerful new tool in the field of object segmentation in the wild. Field datasets are often smaller, and can help encourage discoveries in few-shot learning or simulated data generation. A promising direction for future work includes leveraging synthetic data to augment real sonar datasets for learning-based detection and segmentation tasks (Lee et al., 2018; Sethuraman et al., 2022). Additionally, AI4Shipwrecks narrows a focus on advancing network architecture to enable deep neural networks to learn from limited training data. Of notable interest is recent work that has demonstrated the potential for few shot learning for object detection from marine optical and sonar imagery (Ochal et al., 2020). Few shot and one shot learning approaches aim to effectively learn to represent a class of objects after seeing a few or single instance of that class, which is ideal for datasets with relatively low abundance of samples per class. We hope the AI4Shipwrecks dataset will enable future work on training from limited training data for marine applications.
@@
We would like to acknowledge the lives that were lost in shipwrecks throughout Thunder Bay National Marine Sanctuary. We would also like to thank the sanctuary scientists at Thunder Bay National Marine Sanctuary for supporting field experiments and labeling efforts. Additionally, we acknowledge the effort of IJRR reviewers in providing feedback to this pre-print. The link to the official publication is: https://journals.sagepub.com/doi/10.1177/02783649241266853.